# Unified Batch Normalization: Identifying and Alleviating the Feature Condensation in Batch Normalization and a Unified Framework

### 27 Nov 2023 v1

@@ Abstract
Batch Normalization (BN) has become an essential technique in contemporary neural network design, enhancing training stability. Specifically, BN employs centering and scaling operations to standardize features along the batch dimension and uses an affine transformation to recover features. Although standard BN has shown its capability to improve deep neural network training and convergence, it still exhibits inherent limitations in certain cases. Most existing techniques that enhance BN consider a single or a few aspects of BN. In this paper, we first identify problems with BN from a feature perspective and explore that feature condensation exists in the learning when employing BN, which negatively affects the testing performance. To tackle this problem, we propose a two-stage unified framework called Unified Batch Normalization (UBN). In the first stage, we utilize a simple feature condensation threshold to alleviate the feature condensation, which hinders inappropriate statistic updates in normalization. In the second stage, we unify various normalization variants to boost each component of BN. Our experimental results reveal that UBN significantly enhances performance across different visual backbones and notably expedites network training convergence, particularly in early training stages. Notably, our method improved about 3% in top-1 accuracy on ImageNet classification with large batch sizes, showing the effectiveness of our approach in real-world scenarios.
@@ 1 Introduction
Figure 1: The feature condensation phenomenon and learning dynamics when performing normalization methods on the CIFAR-10 dataset (a) The average feature cosine similarity between input features of the first normalization layers in ResNet-34 [13]. UBN reduces the feature condensation phenomenon by leveraging the running statistics, improving normalization performance. (b) The training loss of ResNet-34 of different representative normalization methods. (c) The testing accuracy of ResNet-34 of different representative normalization methods. UBN shows faster convergence and better performance than BN and other state-of-the-art (SOTA) normalization methods. We used non-uniform axis scale to visualize the differences between UBN and other SOTA normalization methods in terms of loss and accuracy.
Table 1: Study on the rectifications of the three components widely used in BN. Representative normalization variants only address a single component of BN. In comparison, our proposed framework UBN unifies diverse normalization variants and alleviates feature condensation to perform rectifications in all aspects of BN, thereby enhancing the testing performance and training efficiency.
Component BN [16] BNET [35] IEBN [20] RBN [11] UBN
@@
However, in recent years, shortcomings of the BN have been witnessed in some specific applications [32, 2, 30, 24, 9]. Table 1 shows some representative normalization works. Some variants of BN focus on normalization rectification, which involves centering or scaling rectification. For centering rectification, most papers focus on rectifying the input features, i.e., partitioning input features into different groups or different dimensions [31, 23, 33, 18], which may not fully utilize batch statistics. For scaling rectification, [11] tackled the intensity of scaling with additional modification of features, e.g., by leveraging multi-step statistics, or multi-batch statistics [12, 37]. Besides, [35] explores how to improve affine transformation with local channel statistics. In fact, these variants predominantly concentrate on individual or a few aspects within BN without offering a comprehensive solution to rectify the inconsistency in statistics across all components. This limitation has resulted in restricted practical applicability in real-world applications.
To address these challenges, we introduce a comprehensive two-stage method, Unified Batch Normalization (UBN). We first examine BN from a feature learning perspective. We discover a feature condensation phenomenon in DNNs employing BN, which hinders the learning process. We attribute the shortcomings of BN to its failure to alleviate feature condensation effectively. In order to mitigate this issue, in the first stage, we propose a straightforward yet powerful solution termed the “Feature Condensation Threshold,” which determines the running statistics used during training. In the second stage, we establish a unified framework to boost BN, which introduces rectifications for each component of BN to further enhance its performance.
In this work, we summarize the three key components widely used in BN, as shown in Table 1. Our proposed unified framework, UBN, collectively enhances the efficacy of existing normalization methods in deep neural networks. Specifically, contributions can be summarized as follows:
  • We discover a feature condensation phenomenon, which can impede the learning of DNNs with BN.
  • To tackle the feature condensation, we propose a conditional update scheme decided by a threshold to determine the statistics used in normalization.
@@
  • We propose a unified framework that unifies all rectifications of components in BN.
  • We empirically validated the performance of our proposed UBN on renowned image datasets. Each rectification of UBN is validated to be useful for enhancing testing performance. We demonstrate substantial improvements in testing accuracy, outperforming existing variants. Besides, our approach significantly boosts early-stage training efficiency in modern neural network architectures.
@@ 2.1 Normalization Methods
Component-Specific Approaches to Boost BN. Numerous studies have delved into specific components of BN, aiming to enhance or rectify individual aspects such as centering, scaling, and affine operations. Layer Normalization (LN) [1], Group Normalization (GN) [33], Instance Normalization (IN) [31], and Positional Normalization [18] perform different ways of partitioning the input features before centering to leverage statistics from other dimension/position/groups. Representative Batch Normalization (RBN) [11] focuses on centering and scaling adjustments to capture the representative features. BNET [35] rectified the affine operation into a convolutional operation in CNNs to combine local channel information. Recent endeavors have also attempted to amalgamate multiple normalization approaches [22, 26]. Nonetheless, these methodologies predominantly focus on singular or limited aspects of BN, not holistically enhancing its overall performance. Contrarily, our approach presents an integrated framework that amalgamates the best practices in normalization, offering a more comprehensive enhancement to BN’s efficacy.
Running Statistics in BN across iterations. Another focal point in BN research is optimizing running statistics over multiple batches and iterations. Considering the difference between training and inference, several studies have scrutinized the challenges associated with accumulated running statistics. Batch Re-Normalization [15] endeavored to enable the functionality of BN during training with small mini-batches. Memorized Batch Normalization [12] took into account information from multiple recent batches to derive more precise and resilient statistics. AdaBN [19] dynamically adjusted the statistics across all Batch Normalization layers, aiming to achieve profound adaptation effects in the context of domain adaptation tasks. Cross-Iteration Batch Normalization [37] aggregates statistics across multiple training iterations to alleviate the mini-batch dependency. However, these approaches primarily consider neighboring statistics (such as neighboring batches or iterations), potentially neglecting a wider perspective of global running statistics. In contrast, our methodology proposes a novel approach to manage BN’s global statistics by modulating feature condensation, thus offering a more versatile application across various normalization variants.
@@ 2.2 Feature Learning
In the context of feature learning, the examination of feature cosine similarity has played a pivotal role in understanding the feature condensation phenomenon [^1] In this paper, we refer to feature condensation as the phenomenon where the given features have a very high cosine similarity in average., where features exhibit high cosine similarity. This research area has motivated our work, which centers on determining the running statistics in learning by leveraging feature condensation thresholds computed through feature cosine similarity.
Prior research in this domain has yielded valuable insights and techniques. [27] introduced disentangled feature learning techniques that can address the challenge of feature condensation. On the performance front, Cosine Normalization [21] and works on Contrastive Learning [5, 4] have harnessed feature cosine similarity to reduce redundancy and enhance feature representations. [39] proposed dataset condensation to replace the original large training set with a significantly smaller learned set without neglecting importance statistics. [41] tried to explain the condensation of weights in CNNs in the initial learning phase. [40] claimed that BN blocks the first- and second-derivative w.r.t loss functions and affects the feature-learning process in a negative way. Nevertheless, it is worth noting that while these methods have made strides in analyzing dynamics from a condensation perspective, they may need to explicitly tackle the challenge of feature condensation within BN. This limitation motivates our work to explore novel strategies for incorporating feature cosine similarity into BN to mitigate the effects of feature condensation, potentially improving convergence and overall model performance.
Figure 2: The feature condensation heatmap of ResNet-50 on the CIFAR-100 dataset. We randomly sampled 32 images as input and calculated feature cosine similarity of the input features of the first normalization layer of ResNet-50 before, during, and at the end of training. (a) When leveraging BN, the feature condensation became significant during training. (b) When leveraging UBN, the feature condensation phenomenon could be dramatically alleviated compared to BN. Note that the weight initialization of each model randomly decides the feature in the first epoch.
@@ 3 Method
Figure 3: Schematic diagram of the standard BN and our proposed UBN. (a) BN consists of three components, i.e., centering, scaling, and affine transformation. (b) UBN improves BN with a two-stage method. In Stage 1, UBN defines the statistics for computing normalization by comparing the condensation score $S$ of the input features $\mathbf{X}$ with a given feature condensation threshold $\tau$. In Stage 2, UBN performs rectifications on each component of BN with the given statistics determined in Stage 1.
@@ 3.1 Revisiting Batch Normalization
3.2 Identifying the Feature Condensation Phenomenon during training with BN
Previous work mainly focused on improving a single or few components of BN. In contrast, we focus on a more fundamental problem in BN: how to determine the statistics in centering and scaling operations? We aim to analyze it from a feature perspective. We assume that simply determining the normalization statistics according to the training or inference phase is inappropriate in BN because it ignores the difference between samples in each batch. Therefore, we aim to investigate the relationship between the input feature of BN and its relationship with performance.
We first explore the learning dynamics with BN. As Figure 2 shows, the feature similarity is highly reduced when applying UBN in DNNs compared with BN. We investigate the feature condensation in the learning of DNNs with BN. Such feature condensation is closely related to the learning dynamics. When the feature condensation remains to be high, the learning gets slow and eventually causes a lower testing performance. Therefore, we attribute that BN may fail when the feature condensation phenomenon gets severe. To tackle this problem, we design a simple yet effective threshold to determine the statistics used in normalization, which is detailed in Section 3.3.1.
3.3 Unified Batch Normalization
We aim to determine the appropriate statistics for normalization, enhance the reliable statistics used in centering and scaling, improve the recovery steps of affine transformation, and maintain the benefits of BN. Our proposed Unified Batch Normalization (UBN) strengthens batch statistics of specified features and provides performance gain and reliability in training. UBN also exhibits a high efficiency for training DNNs, especially in the initial stage of training.
@@ 3.3.1 Feature Condensation Rectification
Previously, we attributed one main drawback of BN to its failure to effectively alleviate the feature condensation problem. We propose a simple yet effective threshold called Feature Condensation Threshold (FCT) to alleviate the feature condensation phenomenon better. According to Eq. (2), BN uses different mean and variance during training and inference. However, different batches of features may share different distributions. In some cases, the feature similarity in a batch would be high; therefore, using centering and scaling operations would benefit the learning by reducing the similarity between instances; by contrast, when the feature similarity is already low, using the accumulated running mean and variance would stabilize the training.
@@
7:Stage 2: Normalization with given statistics
8: Centering Rectification: $\mathbf{X^{\prime}}=\mathbf{X}+w_{c}\odot\mathbf{K(\mathbf{X})}$
9: Centering: $\mathbf{\bar{X}^{\prime}}=\mathbf{X^{\prime}}-\mu$
10: Scaling: $\mathbf{\hat{X}^{\prime}}=\frac{\mathbf{\bar{X}}^{\prime}}{\sqrt{\sigma^{2}+\varepsilon}}$
11: Scaling Rectification: $\mathbf{\hat{X}^{\prime\prime}}=\mathbf{\hat{X}^{\prime}}\cdot\text{sigmoid}(w_{s}\odot\mathbf{K(\mathbf{\hat{X}^{\prime}})}+b_{s})$
12: Affine Rectification: $\mathbf{Y}=\text{conv}(\mathbf{\hat{X}^{\prime\prime}})$
$S_{B}=\frac{1}{B}\sum_{i,j\in B,i\neq j}\cos(X_{i},X_{j})$ (3)
$S\leftarrow m^{\prime}S+(1-m^{\prime})S_{B}$
@@ 3.3.2 Normalization Rectification
The rectification process involves the introduction of additional trainable weights by adopting the idea from RBN [11]. Specifically, during the centering operation, we learn weights $w_{c}$ to capture the importance of representative features. We apply average pooling on the given input feature and use $w_{c}$ to obtain the residual rectification of our input features. During the scaling operation, we learn weights $w_{s}$ and $b_{s}$ to capture the intensity of de-centered features to maintain better feature intensity. We also apply average pooling on the given decentralized features and use $w_{s}$ and $b_{s}$ to obtain the rectified intensity scaling factor. Specifically, the normalization rectification is shown as follows:
@@ 3.3.3 Affine Rectification
where $\mathbf{Y}$ is obtained using convolutional operation on the channel information of given batch features $\mathbf{X}$.
@@ 4 Experiments
We evaluate UBN in image classification on the ImageNet [25], CIFAR-100 [17], and CIFAR-10 [17] datasets. To validate the design of UBN, we further perform ablation studies on each rectification in each component of BN. Note that UBN is not limited to these evaluations and can benefit a general class of architectures with BN.
4.1 Implementation Details
Table 2: Classification performance of utilizing UBN on multiple network architectures on the ImageNet dataset.
@@ 4.1 Implementation Details
Table 3: Classification performance of utilizing UBN on multiple network architectures on the CIFAR-10 dataset.
CIFAR-10 Normalization Top-1 Accuracy (%) Top-5 Accuracy (%)
ResNet-18 BN 93.65 99.85
UBN($\tau=0.025$) 93.80 99.81
ResNet-34 BN 93.02 99.76
UBN($\tau=0.15$) 93.98 99.83
ResNet-50 BN 92.26 99.84
UBN($\tau=0.15$) 94.32 99.85
ResNet-101 BN 93.19 99.87
UBN($\tau=0.25$) 94.56 99.89
VGG-11 BN 91.14 99.65
UBN($\tau=0.25$) 91.48 99.66
Table 4: Classification performance of utilizing UBN on multiple network architectures on the CIFAR-100 dataset.

### 08 Mar 2024 v2

@@ Abstract
Batch Normalization (BN) has become an essential technique in contemporary neural network design, enhancing training stability. Specifically, BN employs centering and scaling operations to standardize features along the batch dimension and uses an affine transformation to recover features. Although standard BN has shown its capability to improve deep neural network training and convergence, it still exhibits inherent limitations in certain cases. Current enhancements to BN typically address only isolated aspects of its mechanism. In this work, we critically examine BN from a feature perspective, identifying feature condensation during BN as a detrimental factor to test performance. To tackle this problem, we propose a two-stage unified framework called Unified Batch Normalization (UBN). In the first stage, we employ a straightforward feature condensation threshold to mitigate condensation effects, thereby preventing improper updates of statistical norms. In the second stage, we unify various normalization variants to boost each component of BN. Our experimental results reveal that UBN significantly enhances performance across different visual backbones and different vision tasks, and notably expedites network training convergence, particularly in early training stages. Notably, our method improved about 3% in accuracy on ImageNet classification and 4% in mean average precision on both Object Detection and Instance Segmentation on COCO dataset, showing the effectiveness of our approach in real-world scenarios.
Keywords:
Feature Condensation Batch Normalization Image Classification Object Detection Instance Segmentation
@@ 1 Introduction
Table 1: Study on the rectifications of the three components widely used in BN. Classical normalization variants only address a single component of BN. In comparison, our proposed UBN unifies diverse normalization variants and alleviates feature condensation to perform rectifications in all aspects of BN, thereby enhancing the testing performance and training efficiency.
Component BN [16] BNET [37] IEBN [20] RBN [11] UBN (ours)
@@
However, in recent years, shortcomings of the BN have been witnessed in some specific applications [34, 2, 32, 26, 9]. Table 1 shows some representative normalization works. Some variants of BN focus on normalization rectification, which involves centering or scaling rectification. For centering rectification, most papers focus on rectifying the input features, i.e., partitioning input features into different groups or different dimensions [33, 25, 35, 18], which may not fully utilize batch statistics. For scaling rectification, [11] tackled the intensity of scaling with additional modification of features, e.g., by leveraging multi-step statistics, or multi-batch statistics [12, 39]. Besides, [37] explores how to improve affine transformation with local channel statistics. In fact, these variants predominantly concentrate on individual or a few aspects within BN without offering a comprehensive solution to rectify the inconsistency in statistics across all components. This limitation has resulted in restricted practical applicability.
To overcome these challenges, we propose a comprehensive, two-stage approach named Unified Batch Normalization (UBN). Our initial analysis of BN from a feature learning perspective uncovers a feature condensation phenomenon within DNNs utilizing BN, which poses a barrier to effective learning. We ascribe BN’s limitations to its ineffectiveness in counteracting feature condensation. In response, our first stage introduces a simple yet impactful “Feature Condensation Threshold,” guiding the selection of statistics during training. The second stage presents a unified framework to enhance BN, incorporating rectifications across its components to improve overall performance. Our contributions are as follows:
  1. 1.
    To our best knowledge, we are the first to identify the feature condensation phenomenon in BN, a critical insight that sheds light on previously unaddressed impediments to the learning of normalization methods.
  2. 2.
    We introduce an innovative conditional update strategy with the introduction of feature condensation threshold. This approach selectively navigates the normalization statistics during training, effectively mitigating the feature condensation issue and promoting more dynamic and adaptable learning pathways.
  3. 3.
@@
    We propose a unified framework that unifies typical rectifications of components in BN [11, 37, 20], which can be incorporated into different training scenarios.
  4. 4.
    We evaluate UBN’s effectiveness across different image datasets and vision tasks, showcasing notable testing accuracy improvements (e.g., exceeding 3% accuracy gain on ImageNet) and early-stage training efficiency boosts. UBN also achieves significant average precision improvements (e.g., exceeding 4% mean average precision gains on COCO) in both instance segmentation and object detection tasks.
Figure 1: The feature condensation phenomenon and learning dynamics when performing normalization methods on the CIFAR-10 dataset (a) The average feature cosine similarity between input features of the first normalization layers in ResNet-34 [13]. UBN reduces the feature condensation phenomenon by properly leveraging the running statistics, improving normalization performance. (b) The training loss of ResNet-34 of different representative normalization methods. (c) The testing accuracy of ResNet-34 of different representative normalization methods. UBN shows faster convergence and better performance than BN and other state-of-the-art (SOTA) normalization methods. We used non-uniform axis scale to visualize the differences between UBN and other SOTA normalization methods in terms of loss and accuracy.
@@ 2.1 Normalization Methods
2.1.1 Component-Specific Approaches to Boost BN
Numerous studies have delved into specific components of BN, aiming to enhance or rectify individual aspects such as centering, scaling, and affine operations. Layer Normalization (LN) [1], Group Normalization (GN) [35], Instance Normalization (IN) [33], and Positional Normalization [18] perform different ways of partitioning the input features before centering to leverage statistics from other dimension/position/groups. Representative Batch Normalization (RBN) [11] focuses on centering and scaling adjustments to capture the representative features. BNET [37] rectified the affine operation into a convolutional operation in CNNs to combine local channel information. Recent endeavors have also attempted to amalgamate multiple normalization approaches [24, 28]. Nonetheless, these methodologies predominantly focus on singular or limited aspects of BN, not holistically enhancing its overall performance. Contrarily, our approach presents an integrated framework that amalgamates the best practices in normalization, offering a more comprehensive enhancement to BN’s efficacy.
2.1.2 Running Statistics in BN across iterations
Another focal point in BN research is optimizing running statistics over multiple batches and iterations. Considering the difference between training and inference, several studies have scrutinized the challenges associated with accumulated running statistics. Batch Re-Normalization [15] endeavored to enable the functionality of BN during training with small mini-batches. Memorized Batch Normalization [12] took into account information from multiple recent batches to derive more precise and resilient statistics. AdaBN [19] dynamically adjusted the statistics across all Batch Normalization layers, aiming to achieve profound adaptation effects in the context of domain adaptation tasks. Cross-Iteration Batch Normalization [39] aggregates statistics across multiple training iterations to alleviate the mini-batch dependency. However, these approaches primarily consider neighboring statistics (such as neighboring batches or iterations), potentially neglecting a wider perspective of global running statistics. In contrast, our methodology proposes a novel approach to manage BN’s global statistics by modulating feature condensation, thus offering a more versatile application across various normalization variants.
@@ 2.2 Feature Learning
In the context of feature learning, the examination of feature cosine similarity has played a pivotal role in understanding the feature condensation phenomenon, where features exhibit high cosine similarity. This research area has motivated our work, which centers on determining the running statistics in learning by leveraging feature condensation thresholds computed through feature cosine similarity.
Prior research in this domain has yielded valuable insights and techniques. [29] introduced disentangled feature learning techniques that can address the challenge of feature condensation. On the performance front, Cosine Normalization [23] and works on Contrastive Learning [5, 4] have harnessed feature cosine similarity to reduce redundancy and enhance feature representations. [41] proposed dataset condensation to replace the original large training set with a significantly smaller learned set without neglecting importance statistics. [22] discussed a possible reason for the decrease of the feature diversity in the early learning of MLPs without batch normalization. [42] tried to explain the condensation of weights in CNNs in the initial learning phase. [43] claimed that BN blocks the first and second derivatives w.r.t loss functions and affects the feature-learning process in a negative way. Nevertheless, it is worth noting that while these methods have made strides in analyzing dynamics from a condensation perspective, they may need to explicitly tackle the challenge of feature condensation within BN. This limitation motivates our work to explore novel strategies for incorporating feature cosine similarity into BN to mitigate the effects of feature condensation, potentially improving convergence and overall model performance.
Figure 2: The feature condensation heatmap of ResNet-50 on the CIFAR-100 dataset. We randomly sampled 32 images as input and calculated the average cosine similarity of the input features of the first normalization layer of ResNet-50 before, during, and at the end of training. (a) When leveraging BN, the feature condensation became significant during training. (b) When leveraging UBN, the feature condensation phenomenon could be dramatically alleviated compared to BN. Note that the weight initialization of each model randomly decides the feature in the first epoch.
@@ 3.1 Revisiting Batch Normalization
3.2 Identifying the Feature Condensation Phenomenon in Training with BN
Previous work mainly focused on improving a single or few components of BN. In contrast, we focus on a more fundamental problem in BN: how to determine the statistics in centering and scaling operations? We aim to analyze it from a feature perspective. We assume that simply determining the normalization statistics according to the training or inference phase is inappropriate in BN because it ignores the difference between samples in each batch. Therefore, we aim to investigate the relationship between the input feature of BN and its relationship with performance. We formally define the feature condensation phenomenon in BN as follows:
Definition 1 (Feature Condensation)
In high-dimensional feature spaces, feature condensation is identified when a set of features $X$ demonstrates an average cosine similarity $\mathbb{E}_{x,x^{\prime}\in X,x\neq x^{\prime}}\left[\cos(x,x^{\prime})\right]$ exceeding a threshold $\tau$.
We first explore the learning dynamics with BN, and observe a notable pattern as depicted in Figure 2. Specifically, feature similarity remains significantly elevated under BN. This phenomenon, termed feature condensation (Definition 1), plays a pivotal role in the learning process. We observe that when the feature condensation remains to be high, the learning gets slow and eventually causes a lower testing performance.
The underlying cause of BN-induced feature condensation appears to be its normalization strategy, which centralizes means and minimizes variance among feature vectors $x_{i}$ and $x_{j}$. This approach tends to elevate their dot product in relation to their norms, a process stabilized by BN. Hence, this standardized operation increases the average cosine similarity among feature vectors post-BN, often surpassing the threshold $\tau$ and leading to feature condensation as defined. To tackle this problem, we adopt a simple yet effective threshold to set the statistics for normalization, as detailed in Section 3.3.1.
3.3 Alleviating Feature Condensation with Rectifications
Motivated by the limitations of Batch Normalization in scenarios characterized by feature condensation, our goal is to identify the most effective statistics for normalization. We aim to enhance the precision of the statistics used for centering and scaling, improve the affine transformation recovery process, and retain the core advantages of BN. Our proposed solution, Unified Batch Normalization (UBN), specifically improves the batch statistics for targeted features, leading to gains in performance and increased reliability during the training of DNNs.
@@ 3.3.1 Feature Condensation Rectification
Previously, we attributed one main drawback of BN to its failure to effectively alleviate the feature condensation problem. We propose a simple yet effective threshold called Feature Condensation Threshold (FCT) to alleviate the feature condensation phenomenon better. According to Eq. (2), BN uses different mean and variance during training and inference. However, different batches of features may share different distributions. In some cases, the feature similarity in a batch would be high; therefore, using centering and scaling operations would benefit the learning by reducing the similarity between instances; by contrast, when the feature similarity is already low, using the accumulated running mean and variance would stabilize training.
Figure 3: Sketch of the standard BN and our UBN. (a) BN consists of three components, i.e., centering, scaling, and affine transformation. (b) UBN improves BN with a two-stage method. In Stage 1, UBN defines the statistics for computing normalization by comparing the condensation score $S$ of the input features $\mathbf{X}$ with a given feature condensation threshold $\tau$. In Stage 2, UBN performs rectifications on each component of BN with the given statistics determined in Stage 1.
$S_{B}=\frac{1}{B}\sum_{i,j\in B,i\neq j}\cos(X_{i},X_{j})$ (3)
$S\leftarrow m^{\prime}S+(1-m^{\prime})S_{B}$
@@
[…]
8:Stage 2: Normalization with given statistics
9: Centering Rectification: $\mathbf{X^{\prime}}=\mathbf{X}+w_{c}\odot\mathbf{K(\mathbf{X})}$
10: Centering: $\mathbf{\bar{X}^{\prime}}=\mathbf{X^{\prime}}-\mu$
11: Scaling: $\mathbf{\hat{X}^{\prime}}=\frac{\mathbf{\bar{X}}^{\prime}}{\sqrt{\sigma^{2}+\varepsilon}}$
12: Scaling Rectification: $\mathbf{\hat{X}^{\prime\prime}}=\mathbf{\hat{X}^{\prime}}\cdot\text{sigmoid}(w_{s}\odot\mathbf{K(\mathbf{\hat{X}^{\prime}})}+b_{s})$
13: Affine Rectification: $\mathbf{Y}=\text{conv}(\mathbf{\hat{X}^{\prime\prime}})$
@@ 3.3.2 Normalization Rectification
The rectification process involves additional trainable weights by adopting the idea from RBN [11]. Specifically, during the centering operation, we learn weights $w_{c}$ to capture the importance of representative features. We apply average pooling on the given input feature and use $w_{c}$ to obtain the residual rectification of our input features. During the scaling operation, we learn weights $w_{s}$ and $b_{s}$ to capture the intensity of de-centered features to maintain better feature intensity. We also apply average pooling on the given decentralized features and use $w_{s}$ and $b_{s}$ to obtain the rectified intensity scaling factor. Specifically, the normalization rectification is shown as follows:
@@ 3.3.3 Affine Rectification
where $\mathbf{Y}$ is obtained using convolutional operation on the channel of features $\mathbf{X}$.
@@ 4 Experiments
We evaluate UBN in image classification on the ImageNet [27], CIFAR-100 [17], and CIFAR-10 [17] datasets. We also generalize UBN to Object Detection and Instance Segmentation on the COCO [21] dataset. To validate the design of UBN, we further perform ablation studies on each rectification in each component of BN. Note that UBN is not limited to these evaluations and can benefit a general class of architectures.
Table 2: Classification performance of different normalization methods on the ImageNet.
@@
4.1 Implementation Details
4.1.1 On ImageNet
Consistent with common practices, we applied data augmentation and randomly cropped images to $224\times 224$ pixels for training ResNets [13], ResNeXt-50 [36], and Res2Net-50 [10]. ResNet-50 was trained using four Nvidia A100 GPUs and the other networks were trained using eight Nvidia A100 GPUs. Each model was trained for 90 epochs. We adopted a multi-step decay process for the learning rate. The optimizer was SGD, with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 1e-4. The learning rate was programmed to decrease by 0.1 at the 30th and 60th epochs. The batch size per GPU was set to 256. In each model, the BN layer was substituted with our as UBN, detailed in Table 2.
4.1.2 On CIFAR-100
We also conducted experiments on the CIFAR-100 using three Nvidia 3090 GPUs. Our process followed typical data augmentation practices, including random cropping of images to $32$ pixels for training ResNets [13], ResNeXts [36], VGGs [30], and Inception-v4s [31]. Each model underwent 200 training epochs. We introduced a warm-up phase in the first epoch, where the learning rate increased linearly with each batch to stabilize early training. After this, we employed a multi-step decay for the learning rate. Using SGD as the optimizer, we started with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 5e-4. The learning rate was set to decrease by 0.2 at the 60th, 120th, and 160th epochs. We used a batch size of 128 per GPU. The BN layer in each model was replaced with our UBN, as shown in Table 3.
Table 3: Classification performance of different normalization methods on the CIFAR-100.